Please help me solve this equation:
Write the equation of the line passing through each of the given points with the indicated slope. Give your results in slope-intercept form, where possible.
(1, -4), m = -4



Answer by nerdybill(7384)

Write the equation of the line passing through each of the given points with the indicated slope. Give your results in slope-intercept form, where possible.
(1, -4), m = -4
.
Plug given info into the the "point-slope" form of:
y-y1 = m(x-x1)
y-(-4) = (-4)(x-1)
y + 4 = (-4)x + 4
y = -4x (this is what they're looking for)
or you could include the y-intercept:
y = -4x + 0